OpenStudy (anonymous):

How do you rationalize negative square root 3/5?

OpenStudy (eujc21):

sqrt(-1) = i

OpenStudy (anonymous):

Is that the answer?

OpenStudy (eujc21):

no I'm letting you know that \[\sqrt{-1* 3/5} =.....\]

OpenStudy (anonymous):

Oh I was wondering....

OpenStudy (anonymous):

My calculator doesn't do negative uner the square root it says error

OpenStudy (eujc21):

I understand, but this is to fundamentally understand that a -1 under a square root is = to i

OpenStudy (anonymous):

Oh I know that but the negative is outside of the square root

OpenStudy (anonymous):

I learned I last year

OpenStudy (anonymous):

definition: \[\sqrt{(-1)} = i \] using that definition, you should have something like this \[\sqrt{(-3/5)} = \sqrt{(-1)}\sqrt{(3/5)}\] applying the definition you get \[i \sqrt{(3/5)}\]